A taxi company charges riders a fixed charge of $1.60 plus $1.50 per mile. How many miles must a rider go to have an average cost per mile of $1.90?
Answer by rfer(16322)

1.60+1.50x=1.90x
1.60=1.90x-1.50x
1.60=0.4x
1.60/0.40=x
x=4 miles
----------------
1.60+1.50*4=1.90*4
7.60=7.60